# A Kaczmarz-inspired approach to accelerate the optimization of neural network wavefunctions

### v1

@@ 1 Introduction
Another approach to the electronic structure problem is the variational Monte Carlo method (VMC) foulkes2001quantum ; becca2017quantum . In VMC, the properties of the wavefunction are calculated using Monte Carlo sampling rather than direct numerical integration, and the energy is variationally minimized through a stochastic optimization procedure. This increases the cost of the calculations, especially when high accuracy is required, but it enables the use of much more general ansatzes. Traditionally, these ansatzes included Slater-Jastrow wavefunctions and Slater-Jastrow-backflow wavefunctions foulkes2001quantum .
In recent decades the modeling of very high-dimensional data and functions has seen impressive progress through the use of neural networks. For the quantum chemistry problem, the high-dimensional quantum wavefunction can be modeled by combining several neural network layers with a determinantal layer that enforces the Fermionic antisymmetry pfau2020ab ; hermann2020deep . Such neural network wavefunctions, optimized using the variational Monte Carlo method, have enabled near-exact computation of the ground-state energy for small molecules schatzle2023deepqmc ; von2022self and certain condensed matter systems Casella2023 ; kim2023neuralnetwork ; lou2023neural ; pescia2023messagepassing ; li2022ab .
Due to the large number of parameters and the highly nonlinear nature of neural networks, the optimization of neural network wavefunctions poses a significant challenge. Prior to neural network wavefunctions, VMC simulations typically relied on powerful domain-specific optimizers such as stochastic reconfiguration (SR) sorella_sr ; nightingale2001optimization ; sorella2007weak and the linear method umrigar2007 ; toulouse2007optimization ; neuscamman2012optimizing ; zhao2017blocked ; sabzevari2020accelerated . These optimizers are able to converge in only tens or hundreds of iterations using highly accurate gradient estimates based on millions or more Monte Carlo samples. However, when applied to a wavefunction with $N_{p}$ parameters, they require solving either a linear system or a generalized eigenvalue problem involving a dense $N_{p}\times N_{p}$ matrix known as the $S$ matrix, with a cost that is $O(N_{p}^{3})$ in general.
@@
For the FermiNet and related architectures pfau2020ab ; von2022self ; li2022ab ; schatzle2023deepqmc , the most popular optimizer is the Kronecker-Factored Approximate Curvature method (KFAC) martens2015optimizing . KFAC was originally designed as a tractable approximation to natural gradient descent (NGD) amari1998natural for machine learning models with millions of parameters. In the quantum setting, natural gradient descent is equivalent to stochastic reconfiguration since the $S$ matrix can be viewed as the Fisher information matrix of the normalized probability distribution defined by the wavefunction pfau2020ab . Taking the view of stochastic reconfiguration, the key ingredient of KFAC is an approximate factorization of the $S$ matrix. This factorization enables the efficient inversion of the approximate $S$ matrix and also makes it possible to approximate $S$ based on multiple recent minibatches. This can significantly improve performance in highly stochastic settings where a single minibatch can provide only a very noisy approximation to the true $S$ matrix. It is worth noting that the interpretation of KFAC as an approximate natural gradient method has recently been called into question due to some experiments which show that KFAC performs better than exact natural gradient updates benzing2022gradient . Regardless of the underlying mechanism, KFAC represents the state-of-the-art in optimizing neural network wavefunctions for molecules and solids.
@@
An alternative to KFAC is to take advantage of the fact that the estimate of the $S$ matrix is always low-rank when it is based on only a small set of Monte Carlo samples. This idea was first introduced in the machine learning community in the form of efficient subsampled natural gradient descent ren2019efficient . In the physics community, several recent works have proposed similar methods for the VMC setting minsr ; rende2023simple . These methods make fewer heuristic assumptions than KFAC and are much simpler to describe and implement. However, unlike KFAC, they are limited to estimating the $S$ matrix based on only a single minibatch at a time.
@@
In this work, we improve upon these existing optimizers with a new method that we call Subsampled Projected-Increment Natural Gradient Descent (SPRING). Our method is most directly inspired by the minimum-step stochastic reconfiguration (MinSR) approach of Chen and Heyl minsr . MinSR is based on the observation that the SR parameter update can be formulated as the solution to an overdetermined linear least-squares problem. At each iteration, the Monte Carlo samples provide access to a small set of rows from this least-squares problem, yielding an underdetermined subproblem which has many possible solutions. To make the parameter update unique, Chen and Heyl propose to choose the minimal-norm solution to the sampled subproblem.
@@
We improve upon this scheme by taking inspiration from the randomized Kaczmarz method for solving overdetermined linear least-squares problems strohmer2009randomized . The original randomized Kaczmarz method solves the problem by sampling a single row at a time. The subsequently developed block version of the method (see i.e. needell2014paved ) instead uses a small set of rows at each iteration. In either case, the key idea of the Kaczmarz method is to iteratively project the solution vector onto the hyperplane of solutions to each sampled subproblem. When the system is consistent and the rows are sampled from an appropriate probability distribution, this approach provably converges to the true solution with an error that decays by a constant factor at each iteration.
@@
To demonstrate the effectiveness of SPRING, we apply it to optimize a FermiNet wavefunction for several small atoms and molecules, namely the carbon, nitrogen, and oxygen atoms and two configurations of the N2 molecule. We use the VMCNet code vmcnet2021github for all numerical experiments. We find that SPRING consistently outperforms MinSR and KFAC when all three methods have their learning rates tuned for maximal performance. Remarkably, on the oxygen atom, SPRING is able to attain chemical accuracy after forty thousand training iterations, whereas both MinSR and KFAC fail to do so even after one hundred thousand iterations.
@@ 2.1 Variational Monte Carlo
In this work we focus on the application of the variational Monte Carlo method to find ground-state wavefunctions for molecular systems. A molecule is defined by the positions $R_{I}$ and charges $Z_{I}$ of $M$ atomic nuclei. Using the Born-Oppenheimer approximation born1927 , we model the nuclei as fixed point charges. The quantum system of interest then consists of $N$ electrons interacting under the influence of these charges. The Hamiltonian for the system is given in atomic units by
$H=-\frac{1}{2}\Delta_{R}-\sum_{i=1}^{N}\sum_{I=1}^{M}\frac{Z_{I}}{\left|r_{i}-R_{I}\right|}+\sum_{i<j}^{N}\frac{1}{\left|r_{i}-r_{j}\right|}+\sum_{I<J}^{M}\frac{Z_{I}Z_{J}}{\left|R_{I}-R_{J}\right|}.$ (1)
Because electrons are fermions, the wavefunction must be antisymmetric with respect to particle exchange. Within the space of antisymmetric wavefunctions, the ground-state energy $E_{0}$ and wavefunction $\left|\psi_{0}\right\rangle$ correspond to the smallest eigenvalue of $H$ and its corresponding eigenfunction.
We make several simplifications to the space of wavefunctions before searching for the ground-state. First, since the Hamiltonian is Hermitian, its ground-state wavefunction can be chosen to be purely real. We thus limit our search to real normalizable many-body wavefunctions of the form $\psi:\mathbb{R}^{3N}\rightarrow\mathbb{R}$, usually written as $\psi(R)=\psi(r_{1},\ldots,r_{N})$ with $R\in\mathbb{R}^{3N}$, $r_{i}\in\mathbb{R}^{3}$. Next, we fix the numbers $N_{\uparrow},N_{\downarrow}$ of up- and down-spin electrons a priori for each calculation. Given this assumption, together with the antisymmetry of the wavefunction and the spin-independence of the Hamiltonian, it is possible to assume without loss of generality that the first $N_{\uparrow}$ electrons are always spin-up and the last $N_{\downarrow}$ electrons are always spin-down foulkes2001quantum .
@@
where $p(R)=\psi_{\theta}(R)^{2}/\left\langle\psi_{\theta}\middle|\psi_{\theta}\right\rangle$ and $E_{L}(R)=H\psi_{\theta}(R)/\psi_{\theta}(R)$ (see i.e. foulkes2001quantum , Section III.C). This last term, $E_{L}(R)$, is known as the local energy of the wavefunction $\left|\psi_{\theta}\right\rangle$ at the position $R$.
To optimize a variational wavefunction using this formula, it is usually necessary to calculate the gradient $\nabla_{\theta}L(\theta)$. A convenient formula for this gradient can be derived taking advantage of the fact that $H$ is Hermitian (see i.e. lin2023explicitly , Appendix E):
@@ 2.2 Stochastic Reconfiguration
One popular approach for optimizing variational wavefunctions is known as stochastic reconfiguration (SR) sorella_sr ; nightingale2001optimization ; sorella2007weak . In recent years, SR has been widely adopted for optimizing neural quantum states in second quantization carleo2017solving ; carleo2019netket . It also serves as the starting point for the MinSR method to be introduced later.
@@
In practice, we make a first order approximation to both $\left|\psi^{\prime}\right\rangle$ and $\left|\psi_{\theta+d\theta}\right\rangle$ and then minimize a second-order approximation to $D^{2}(\left|\psi^{\prime}\right\rangle,\left|\psi_{\theta+d\theta}\right\rangle)$. See park2020geometry for a more thorough discussion of the geometric underpinnings of stochastic reconfiguration.
@@
$\bar{\epsilon}(R)=-\delta\tau\left(E_{L}(R)-L(\theta)\right).$ (10)
@@
A detailed derivation of this formulation of stochastic reconfiguration can be found in minsr .
In practice, (11) must always be approximated using a finite collection of $N_{s}$ samples $R_{1},\ldots,R_{N_{s}}$. The estimate of the Fubini-Study distance of the update $d\theta$ can then be written in matrix form as
@@
where
$\bar{O}=\frac{1}{\sqrt{N_{s}}}\begin{bmatrix}\bar{O}(R_{1})\\ \vdots\\ \bar{O}(R_{N_{s}})\end{bmatrix},\;\bar{\epsilon}=\frac{1}{\sqrt{N_{s}}}\begin{bmatrix}\bar{\epsilon}(R_{1})\\ \vdots\\ \bar{\epsilon}(R_{N_{s}})\end{bmatrix}.$ (14)
The update then takes the form
$d\theta=\underset{d\theta^{\prime}}{\operatorname{argmin}}\;\left\lVert\bar{O}d\theta^{\prime}-\bar{\epsilon}\right\rVert^{2}$ (15)
@@
$d\theta=(\bar{O}^{T}\bar{O}+\lambda I)^{-1}\bar{O}^{T}\epsilon.$ (17)
@@
$d\theta=\frac{\delta\tau}{2}(S+\lambda I)^{-1}g,$ (18)
where $S=\bar{O}^{T}\bar{O}$, $g=\nabla_{\theta}L(\theta)=2\bar{O}^{T}\bar{\epsilon}/\delta\tau$, and the imaginary time step has been halved to cancel the factor of two in the gradient.
@@ 2.3 Minimum-Step Stochastic Reconfiguration
For a problem with $N_{p}$ variational parameters, the traditional SR update scales as $O(N_{p}^{3})$ due to the need to invert the $N_{p}\times N_{p}$ matrix $S$. The insight of MinSR is that the parameter update can be calculated much more efficiently when $N_{s}\ll N_{p}$. In fact the linear system is highly underdetermined in this setting. The MinSR method addresses this issue by choosing the solution with minimal norm. For regularization, MinSR utilizes a pseudoinverse with a cutoff for small eigenvalues. Subsequently, Rende et al rende2023simple suggested using a Tikhonov regularization instead. This approach amounts to solving the same regularized equation as in traditional SR, namely (16). In the underdetermined setting, the application of the Sherman-Morrison-Woodbury formula to (17) yields the solution
@@
This is the version of MinSR that we use as the basis for SPRING and in our numerical experiments. The key point is that the update can now be calculated efficiently since $\bar{O}\bar{O}^{T}$ is only $N_{s}\times N_{s}$ rather than $N_{p}\times N_{p}$. It is worth noting that this method is equivalent to the efficient subsampled natural gradient method proposed by Ren and Goldfarb in 2019 ren2019efficient . We elaborate on the connection between these two methods in A.

### v2

@@ 1 Introduction
Another approach to the electronic structure problem is the variational Monte Carlo method (VMC) [1, 2]. In VMC, the properties of the wavefunction are calculated using Monte Carlo sampling rather than direct numerical integration, and the energy is variationally minimized through a stochastic optimization procedure. This increases the cost of the calculations, especially when high accuracy is required, but it enables the use of much more general ansatzes. Traditionally, these ansatzes included Slater-Jastrow wavefunctions and Slater-Jastrow-backflow wavefunctions [1].
In recent decades the modeling of very high-dimensional data and functions has seen impressive progress through the use of neural networks. For the quantum chemistry problem, the high-dimensional quantum wavefunction can be modeled by combining several neural network layers with a determinantal layer that enforces the Fermionic antisymmetry [3, 4]. Such neural network wavefunctions, optimized using the variational Monte Carlo method, have enabled near-exact computation of the ground-state energy for small molecules [5, 6, 7] and certain condensed matter systems [8, 9, 10, 11, 12].
Due to the large number of parameters and the highly nonlinear nature of neural networks, the optimization of neural network wavefunctions poses a significant challenge. Prior to neural network wavefunctions, VMC simulations typically relied on powerful domain-specific optimizers such as stochastic reconfiguration (SR) [13, 14, 15] and the linear method [16, 17, 18, 19, 20]. These optimizers are able to converge in only tens or hundreds of iterations using highly accurate gradient estimates based on millions or more Monte Carlo samples. However, when applied to a wavefunction with $N_{p}$ parameters, they require solving either a linear system or a generalized eigenvalue problem involving a dense $N_{p}\times N_{p}$ matrix known as the $S$ matrix, with a cost that is $O(N_{p}^{3})$ in general. The recently proposed Rayleigh-Gauss-Newton optimizer [21] presents a potentially favorable intermediate point between stochastic reconfiguration and the linear method, but still incurs a similar computational cost.
@@
For the FermiNet and related architectures [3, 6, 12, 5], the most popular optimizer is the Kronecker-Factored Approximate Curvature method (KFAC) [22]. KFAC was originally designed as a tractable approximation to natural gradient descent (NGD) [23] for machine learning models with millions of parameters. In the quantum setting, natural gradient descent is equivalent to stochastic reconfiguration since the $S$ matrix can be viewed as the Fisher information matrix of the normalized probability distribution defined by the wavefunction [3]. Taking the view of stochastic reconfiguration, the key ingredient of KFAC is an approximate factorization of the $S$ matrix. This factorization enables the efficient inversion of the approximate $S$ matrix and also makes it possible to approximate $S$ based on multiple recent minibatches. This can significantly improve performance in highly stochastic settings where a single minibatch can provide only a very noisy approximation to the true $S$ matrix. It is worth noting that the interpretation of KFAC as an approximate natural gradient method has recently been called into question due to some experiments which show that KFAC performs better than exact natural gradient updates [24]. Regardless of the underlying mechanism, KFAC represents the state-of-the-art in optimizing neural network wavefunctions for molecules and solids.
@@
An alternative to KFAC is to take advantage of the fact that the estimate of the $S$ matrix is always low-rank when it is based on only a small set of Monte Carlo samples. This idea was first introduced in the machine learning community in the form of efficient subsampled natural gradient descent [25]. In the physics community, several recent works have proposed similar methods for the VMC setting [26, 27]. These methods make fewer heuristic assumptions than KFAC and are much simpler to describe and implement. However, unlike KFAC, they are limited to estimating the $S$ matrix based on only a single minibatch at a time.
@@
In this work, we improve upon these existing optimizers with a new method that we call Subsampled Projected-Increment Natural Gradient Descent (SPRING). Our method is most directly inspired by the minimum-step stochastic reconfiguration (MinSR) approach of Chen and Heyl [26]. MinSR is based on the observation that the SR parameter update can be formulated as the solution to an overdetermined linear least-squares problem. At each iteration, the Monte Carlo samples provide access to a small set of rows from this least-squares problem, yielding an underdetermined subproblem which has many possible solutions. To make the parameter update unique, Chen and Heyl propose to choose the minimal-norm solution to the sampled subproblem.
@@
We improve upon this scheme by taking inspiration from the randomized Kaczmarz method for solving overdetermined linear least-squares problems [28]. The original randomized Kaczmarz method solves the problem by sampling a single row at a time. The subsequently developed block version of the method (see i.e. [29]) instead uses a small set of rows at each iteration. In either case, the key idea of the Kaczmarz method is to iteratively project the solution vector onto the hyperplane of solutions to each sampled subproblem. When the system is consistent and the rows are sampled from an appropriate probability distribution, this approach provably converges to the true solution with an expected error that decays by a constant factor at each iteration.
@@
To demonstrate the effectiveness of SPRING, we apply it to optimize a FermiNet wavefunction for several small atoms and molecules, namely the carbon, nitrogen, and oxygen atoms, two configurations of the N2 molecule, and the CO molecule at equilibrium. We use the VMCNet code [30] for all numerical experiments. We find that SPRING consistently outperforms MinSR and KFAC when all three methods have their learning rates tuned for maximal performance. Remarkably, on the oxygen atom, SPRING is able to attain chemical accuracy after forty thousand training iterations, whereas both MinSR and KFAC fail to do so even after one hundred thousand iterations.
@@ 2.1 Variational Monte Carlo
In this work we focus on the application of the variational Monte Carlo method to find ground-state wavefunctions for molecular systems. A molecule is defined by the positions $\mathsf{R}_{I}$ and charges $Z_{I}$ of $M$ atomic nuclei. Using the Born-Oppenheimer approximation [31], we model the nuclei as fixed point charges. The quantum system of interest then consists of $N$ electrons interacting under the influence of these charges. The Hamiltonian for the system is given in atomic units by
$H=-\frac{1}{2}\sum_{i=1}^{N}\Delta_{r_{i}}-\sum_{i=1}^{N}\sum_{I=1}^{M}\frac{Z_{I}}{\left|r_{i}-\mathsf{R}_{I}\right|}+\sum_{i<j}^{N}\frac{1}{\left|r_{i}-r_{j}\right|}+\sum_{I<J}^{M}\frac{Z_{I}Z_{J}}{\left|\mathsf{R}_{I}-\mathsf{R}_{J}\right|},$ (1)
where $r_{i}\in\mathbb{R}^{3}$ represents the position of electron $i$ and $\Delta_{r_{i}}$ represents the corresponding Laplacian operator. Because electrons are fermions, the wavefunction must be antisymmetric with respect to particle exchange. Within the space of antisymmetric wavefunctions, the ground-state energy $E_{0}$ and wavefunction $\left|\psi_{0}\right\rangle$ correspond to the smallest eigenvalue of $H$ and its corresponding eigenfunction.
We make several simplifications to the space of wavefunctions before searching for the ground-state. First, since the Hamiltonian is Hermitian, its ground-state wavefunction can be chosen to be purely real. We thus limit our search to real normalizable many-body wavefunctions of the form $\psi:\mathbb{R}^{3N}\rightarrow\mathbb{R}$, usually written as $\psi(R)=\psi(r_{1},\ldots,r_{N})$ with $R\in\mathbb{R}^{3N}$. Next, we fix the numbers $N_{\uparrow},N_{\downarrow}$ of up- and down-spin electrons a priori for each calculation. Given this assumption, together with the antisymmetry of the wavefunction and the spin-independence of the Hamiltonian, it is possible to assume without loss of generality that the first $N_{\uparrow}$ electrons are always spin-up and the last $N_{\downarrow}$ electrons are always spin-down [1].
@@
where $p(R)=\psi_{\theta}(R)^{2}/\left\langle\psi_{\theta}\middle|\psi_{\theta}\right\rangle$ and $E_{L}(R)=H\psi_{\theta}(R)/\psi_{\theta}(R)$ (see i.e. [1], Section III.C). This last term, $E_{L}(R)$, is known as the local energy of the wavefunction $\left|\psi_{\theta}\right\rangle$ at the position $R$.
To optimize a variational wavefunction using this formula, it is usually necessary to calculate the gradient $\nabla_{\theta}L(\theta)$. A convenient formula for this gradient can be derived taking advantage of the fact that $H$ is Hermitian (see i.e. [32], Appendix E):
@@ 2.2 Stochastic Reconfiguration
One popular approach for optimizing variational wavefunctions is known as stochastic reconfiguration (SR) [13, 14, 15] . In recent years, SR has been widely adopted for optimizing neural quantum states in second quantization [33, 34]. It also serves as the starting point for the MinSR method to be introduced later.
@@
In practice, we make a first order approximation to both $\left|\psi^{\prime}\right\rangle$ and $\left|\psi_{\theta+d\theta}\right\rangle$ and then minimize a second-order approximation to $D^{2}(\left|\psi^{\prime}\right\rangle,\left|\psi_{\theta+d\theta}\right\rangle)$. See [35] for a more thorough discussion of the geometric underpinnings of stochastic reconfiguration.
@@
$\bar{\epsilon}(R)=-\delta\tau\left(E_{L}(R)-\mathbb{E}_{R\sim p}\left[E_{L}(R)\right]\right).$ (10)
@@
A detailed derivation of this formulation of stochastic reconfiguration can be found in [26].
In practice, (11) must always be approximated using a finite collection of $N_{s}$ samples $R_{1},\ldots,R_{N_{s}}$. To estimate $\bar{O}(R)$ and $\bar{\epsilon}(R)$ at the sampled points, we must empirically estimate the expectation values that appear in their definitions. We first define
$O=\frac{1}{\sqrt{N_{s}}}\begin{bmatrix}\nabla_{\theta}\log\left|\psi_{\theta}(R_{1})\right|\\ \vdots\\ \nabla_{\theta}\log\left|\psi_{\theta}(R_{N_{S}})\right|\end{bmatrix},\;\epsilon=-\frac{\delta\tau}{\sqrt{N_{s}}}\begin{bmatrix}E_{L}(R_{1})\\ \vdots\\ E_{L}(R_{N_{s}})\end{bmatrix}.$ (13)
Next, let $\mathbf{1}$ be the column vector of length $N_{S}$ whose entries are all equal to $1$ and let $P=\frac{1}{N_{s}}\mathbf{1}\mathbf{1}^{T}$ represent the orthogonal projector onto the span of $\mathbf{1}$. When acting on the left, the operator $P$ replaces each row of the matrix or vector on its right with the average of all of its rows. Our empirical estimates of $\bar{O}(R)$ and $\bar{\epsilon}(R)$ can thus be collected as
$\bar{O}=(I-P)O,\;\bar{\epsilon}=(I-P)\epsilon.$ (14)
The estimate of the Fubini-Study distance of the update $d\theta$ can then be written as
@@
and the update takes the form
$d\theta=\underset{d\theta^{\prime}}{\operatorname{argmin}}\;\left\lVert\bar{O}d\theta^{\prime}-\bar{\epsilon}\right\rVert^{2}.$ (16)
@@
$d\theta=(\bar{O}^{T}\bar{O}+\lambda I)^{-1}\bar{O}^{T}\bar{\epsilon}.$ (18)
@@
$d\theta=-\frac{\delta\tau}{2}(S+\lambda I)^{-1}g,$ (19)
where $S=\bar{O}^{T}\bar{O}$, $g=\nabla_{\theta}L(\theta)=-2\bar{O}^{T}\bar{\epsilon}/\delta\tau$, and the imaginary time step has been halved to cancel the factor of two in the gradient.
@@ 2.3 Minimum-Step Stochastic Reconfiguration
For a problem with $N_{p}$ variational parameters, the traditional SR update scales as $O(N_{p}^{3})$ due to the need to invert the $N_{p}\times N_{p}$ matrix $S$. The insight of MinSR is that the parameter update can be calculated much more efficiently when $N_{s}\ll N_{p}$. In fact the linear system is highly underdetermined in this setting. The MinSR method addresses this issue by choosing the solution with minimal norm. For regularization, MinSR utilizes a pseudoinverse with a cutoff for small eigenvalues. Subsequently, Rende et al [27] suggested using a Tikhonov regularization instead. This approach amounts to solving the same regularized equation as in traditional SR, namely (17). In the underdetermined setting, the application of the Sherman-Morrison-Woodbury formula to (18) yields the solution
@@
This is the version of MinSR that we use as the basis for SPRING. The key point is that the update can now be calculated efficiently since $\bar{O}\bar{O}^{T}$ is only $N_{s}\times N_{s}$ rather than $N_{p}\times N_{p}$.
It is worth noting that methods very similar to MinSR were introduced independently in the machine learning community as early as 2019. For example, Zhang, Martens, and Grosse decribed a pseudoinverse-based approach to natural gradient descent for the overparameterized case when the number of parameters is greater than the total number of data-points available [36] . Subsequently, Ren and Goldfarb proposed an efficient subsampled natural gradient method [25] for the case when the minibatch size is smaller than the number of parameters. In fact, the Tikhonov-regularized form of MinSR can be viewed as a streamlined implementation of the method of Ren and Goldfarb which applies whenever the gradient of the loss function is a linear combination of the model gradients at the sampled points. We elaborate on this connection in A.
2.4 Kaczmarz method for solving linear systems
MinSR enables the efficient solution of an underdetermined subsample from the SR equation, but the solution of such an equation is not necessarily a good approximation to the solution of the original SR equation. Our goal is to develop an algorithm that can converge to the solution of the original SR equation using only a series of underdetermined subsamples. This is the motivation behind the Kaczmarz algorithm, and in particular its blocked variants.
To understand the Kaczmarz algorithm, consider an overdetermined system of linear equations $Ax=b$, where $A$ is $m\times n$ and
$A=\begin{bmatrix}a_{1}^{T}\\ \vdots\\ a_{m}^{T}\end{bmatrix},\;b=\begin{bmatrix}b_{1}\\ \vdots\\ b_{m}\end{bmatrix}.$ (21)
Assume that $A$ has full column rank and that the system is consistent, meaning that there exists a unique $x^{*}$ such that $Ax^{*}=b$. The earliest form of the Kaczmarz algorithm, which dates back to the first half of the twentieth century [37], solves this problem iteratively. A starting guess $x_{0}$ is required for initialization and is then honed by iterations of the form
$x_{k}=x_{k-1}+\frac{a_{i}}{\left\lVert a_{i}\right\rVert^{2}}(b_{i}-a_{i}^{T}x_{k-1}),$ (22)
where $i$ is chosen to cycle through all the rows of $A$ one at a time. The interpretation is that $x_{k}$ is attained by projecting $x_{k-1}$ onto the solution space of the sampled equation $a_{i}^{T}x=b_{i}$.
The seminal work of Strohmer and Vershynin [28] proposed a randomized variant of the algorithm which samples row $i$ with a probability proportional to $\left\lVert a_{i}\right\rVert^{2}$. With this randomized sampling strategy, they showed that the algorithm converges as
$\mathbb{E}\left[\left\lVert x_{k}-x^{*}\right\rVert^{2}\right]\leq\left(1-\kappa_{D}(A)^{-2}\right)^{k}\left\lVert x_{0}-x^{*}\right\rVert^{2},$ (23)
where $\kappa_{D}(A)=\left\lVert A\right\rVert_{F}\left\lVert A^{+}\right\rVert_{2}$ is the Demmel condition number of $A$ and $A^{+}$ is the Moore-Penrose pseudoinverse of $A$.
Since single row updates are very inefficient on modern hardware, blocked versions of the randomized Kaczmarz algorithm were subsequently proposed and analyzed, i.e. [29]. Such randomized block Kaczmarz algorithms take the form
$x_{k}=x_{k-1}+A_{\sigma}^{+}(b_{\sigma}-A_{\sigma}x_{k-1}),$ (24)
where $\sigma$ represents a randomly selected subset of the rows of $A$ and $b$ and $A_{\sigma}^{+}$ represents the pseudoinverse of $A_{\sigma}$. The interpretation is similar to before: $x_{k}$ represents the projection of $x_{k-1}$ onto the solution space of the sampled equation $A_{\sigma}x=b_{\sigma}$. For consistent equations the randomized block Kaczmarz algorithm converges linearly to the true solution $x^{*}$, though the convergence rate is more complicated than in the single row case and has a more nuanced dependence on the sampling procedure.